A car travels 550 mi. A second car, traveling 5 mph faster than the first car, makes the same trip in 1 h less time. Find the speed of each car.
Answer by jorel1380(3719)

550/n=550/(n+5)+1
550n+2750=550n+n2+5n
n2+5n-2750=0
(n+55)(n-50)=0
n=-55 or 50
Throwing out the negative answer, we get the speed of the slower car to be 50 mph, and that of the faster car to be 55 mph..